OpenStudy (swag):

Jason considered two similar televisions at a local electronics store. The generic version was based off the brand name and was three-fourths the size of the brand name. If the generic television set is 12 inches by 24 inches, what are dimensions of the brand name television? Answer 6 inches by 12 inches 9 inches by 18 inches 16 inches by 32 inches 18 inches by 36 inches

OpenStudy (anonymous):

9 inches by 18 inches

OpenStudy (savvy):

16 inches by 32 inches since the size is 3/4th hence alll sides are 3/4th. so (3/4)x=12inches (3/4)y=24inches so x=16, y=32

OpenStudy (swag):

:/

OpenStudy (anonymous):

Oop, I misread the post! b = 4 g/3 => 12 * 4/3 = 16 => 24 * 4/3 = 32